my question is how do I solve the sum of two consecutive even integers is 16 less then three times the larger number find the integers I'm struggling a bit on that question
Answer by rfer(16322)

x+x+2=3(x+2)-16
2x+2=3x+6-16
2x-3x=6-16-2
-x=-12
x=12
x+2=14 largest
-------------------
12+12+2=3(12+2)-16
26=26